Evaluate the following limits,if they exist,where [x] is the greatest interger function


a)lim [2x]/x as x approaches 0


b)lim x[1/x] as x approaches 0


1
Expert's answer
2022-03-28T16:04:26-0400

"a:\\\\x\\in \\left( -\\frac{1}{2},0 \\right) :\\frac{\\left[ 2x \\right]}{x}=\\frac{-1}{x}\\\\\\underset{x\\rightarrow 0-}{\\lim}\\frac{\\left[ 2x \\right]}{x}=-\\infty \\\\x\\in \\left( 0,\\frac{1}{2} \\right) :\\frac{\\left[ 2x \\right]}{x}=0\\\\\\underset{x\\rightarrow 0+}{\\lim}\\frac{\\left[ 2x \\right]}{x}=0\\\\The\\,\\,\\lim it\\,\\,doesn\u0091t\\,\\,exist\\\\b:\\\\x\\cdot \\left( \\frac{1}{x}-1 \\right) <x\\left[ \\frac{1}{x} \\right] \\leqslant x\\cdot \\frac{1}{x}\\\\\\underset{x\\rightarrow 0}{\\lim}x\\left( \\frac{1}{x}-1 \\right) =1,\\underset{x\\rightarrow 0}{\\lim}x\\cdot \\frac{1}{x}=1\\Rightarrow \\\\\\Rightarrow \\underset{x\\rightarrow 0}{\\lim}x\\left[ \\frac{1}{x} \\right] =1"